OpenStudy (anonymous):

If I have 9+5 - (-6) 14- (-6) I know I am supposed to change the signs , but I am not sure how . What would the next process be in solving ?

jimthompson5910 (jim_thompson5910):

Think of "-6" as 6 IOUs. In other words, 6 pieces of paper which represent debt. If you take away those 6 pieces of debt, that's the same as adding $6 to your wallet/purse. So that's why 14 - (-6) = 14 + 6 the two negatives cancel to form a positive

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

So the two negatives become a positive making the answer 20, correct? which would mean if it was opposite and the problem was 14+(-6) then it would basically be 14-6 , right?

jimthompson5910 (jim_thompson5910):

14+(-6) = 14-6, correct

jimthompson5910 (jim_thompson5910):

14+(-6) means "add on 6 bills of debt. So the balance goes from $14 to $8"